OpenStudy (anonymous):

Identify the maximum and minimum values of the function y = 8 cos x in the interval [-2π, 2π].

OpenStudy (whpalmer4):

what is the range of \(\cos x\) over that interval?

OpenStudy (anonymous):

I have no entire clue. I don't understand this at all.

OpenStudy (whpalmer4):

What's the largest value that \(\cos x\) ever has?

OpenStudy (anonymous):

8?

OpenStudy (whpalmer4):

I asked about \(\cos x\), not \(8\cos x\)

OpenStudy (anonymous):

I dont know

OpenStudy (whpalmer4):

here's a hint:

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

right. and the minimum value?

OpenStudy (anonymous):

-1. so for 8 cos x, the max would be 8 & the min would be -8?

OpenStudy (whpalmer4):

you got it!

OpenStudy (anonymous):

so what does it mean by "in the interval [-2pi, 2pi]?

OpenStudy (whpalmer4):

well, in this case, it doesn't really affect the answer. if the interval had been something smaller, where you didn't have an entire cycle of \(\cos x\), the answer might have been different. For example, if the interval was \([0,\pi/4]\) then you would have to find the maximum and minimum values of \(\cos x\) between those two points. The max would still be 1, the min would be \(\sqrt{2}/2\), and the answer to the question would be 8 and \(4\sqrt{2}\) for max and min respectively.

OpenStudy (anonymous):

Okay, thanks so much!

OpenStudy (marissalovescats):

Because the amplitude is 8, then the height of each wave will be 8. So your max would be 8 and your min would be -8. I just had this question in my online class but with the number 10.